1. A McDonald’s Big Mac costs 2.44 yuan in China, but costs $4.20 in the United States. Assuming that purchasing-power parity (PPP) holds, how many Chinese yuan are required to purchase 1 U.S. dollar?
1
Expert's answer
2021-07-30T09:22:02-0400

As per the information given in the question we have :

Cost of McDonald’s Big Mac in China = 2.44 Yuan

Cost of McDonald’s Big Mac in US Dollars = $ 4.20

Since Purchasing power parity holds good, the spot exchange rate between the Yuan and the US Dollar is

"= \frac{Cost\; of \;McDonald’s \;Big\; Mac\; in \;China }{ Cost \;of \;McDonald’s \;Big\; Mac \;in \;US \;Dollars} \\

= \frac{2.44 \;Yuan }{ 4.20 \;Dollar} \\

= 0.5810 \;Yuan"

Thus

"\frac{Yuan }{ Dollar}= 0.5810 \;Yuan"

= 0.58 Yuan ( If rounded off to two decimal places )

Thus when the purchasing power parity holds good, 0.5810 Chinese yuan are required to purchase 1 U.S. dollar.